OpenStudy (anonymous):

Given a = -3, b = 2, h = 1, and k = -4, what is the equation of the graph if the parent function is y=√x? Describe the transformation.

OpenStudy (anonymous):

I don't mind if I only get the answer, but if you can, could you explain it to me please? Thank you >.<

OpenStudy (jdoe0001):

what's the ABH form you have again?

OpenStudy (jdoe0001):

you can just take a screenshot, if it's quicker :)

OpenStudy (anonymous):

y = a√(b(x-h))+k

OpenStudy (anonymous):

Math is a pain the the butt lmao

OpenStudy (jdoe0001):

so A is a shrink or stretch on the graph h is just a horizontal shift k is just the vertical shift

OpenStudy (jdoe0001):

now "b" is a horizontal shift too BUT, the value of B changes the parent function position to h/b h/b > 0, to the left h/b < 0, to the right

OpenStudy (anonymous):

Okay I get that part :)

OpenStudy (jdoe0001):

so the \(\bf \sqrt{x}\) a = 3 stretches the graph I think by 1/3 b =2 and h = 1 h/b = 1/2 positive so the shrunken graph gets moved by 1/2 to the left, that is at x = 0.5 k = -4 the shifted shrunken graph then gets flushed down with a vertical shift to the bottom by 4 units

OpenStudy (jdoe0001):

x = -0.5 I meant, negative that is